# Benchmarking Misuse Mitigation Against Covert Adversaries

### v1

@@ Appendix F Experiments Discussions
In Section 5, we fine-tune Qwen2.5-7B on 700 set of decomposition sub-queries samples collected from o3-mini on MMLU-auxilary, where the number of decompositions varied from 3 to 6 across the samples. The fine-tuning data is highly structured– extensive fine-tuning on such data can lead to overfitting, causing the model to lose its general capabilities. It is known that a fine-tuned model on small or structured datasets memorizes training data rather than overall generalization [71]. Hence, we found that a small learning rate ($61{\rm e}-$ 6) with only 2 epochs work the best to circumvent this. We provide an example question and decomposition datapoint below.
MMLU-auxiliary question Question: Every student in Australia must wear their school uniforms when they are at school. There are many kinds of uniforms such as suits, dresses, shorts, skirts, shirts, T-shirts and so on. Class usually begins at 8:00 a.m. in most Australian schools, and sometimes it begins at 10:00 a.m. You needn’t go to class with your schoolbag, you can put it in your own cabinet and only take a few books. There are many subjects: English, Maths, Biology, Geography, History, Physics, Chemistry, Music, Computer, PE, Chinese Drama and so on. In a computer class, each student sits at a table and does their work. They are very ____. The teacher goes around them to give directions to them. One class is about two hours. At noon all the students have lunch at school. Their lunch is sandwiches and some fruit. In the afternoon there’s only one class. School ends at 3:30 P.m. Some students can swim or play ball games until 5:30 P.m. Then they must go home. They seldom do homework, even the students in Grade Eleven. This is the school life of the Students in Australia. In a computer class, each student is very ____. Choices:
@@
Decomp 1)In the sentence ’In a computer class, each student sits at a table and does their work. They are very ____.’, what setting is being described, and what are the main actions of the students?
Decomp 2)Based on the sentence ’In a computer class, each student sits at a table and does their work. They are very ____.’, is the missing adjective more likely to describe the students’ volume level (how much noise they make) or their emotional state?
Decomp 3)Which of the following words describes a low level of sound: quiet, noisy, happy, or sad?
Decomp 4)What does the phrase ’does their work’ suggest about how much talking or noise the students are making in the computer class context?
Decomp 5)In a scenario where students are each working individually on computers and not talking, would you describe the atmosphere as more ’quiet’ or more ’noisy’?
Decomp 6)Given the sentence ’Each student sits at a table and does their work. They are very ____.’ and the four options quiet, noisy, happy, sad, which word logically fills the blank to match the described atmosphere?
F.1 Decomposition Finetuning Ablation
In order to show that the fine-tuning is targeted to help with the quality of the decomposed sub-queries, and does not significantly enhance the general capability of the fine-tuned model, we provide a baseline experiment where the fine-tuned model is used together with Qwen in the decomposition baseline (i.e., fine-tuned model $\to$ Qwen $\to$ Qwen). We compare this with the scenario that Qwen is used for all three parts. As Table 3 shows, the fine-tuned model does not improve the accuracy when no strong model is deployed.
@@ F.1 Decomposition Finetuning Ablation
Table 3: Distillation leads to misuse because the attacker learns better decomposition strategies, not general gains in capabilities. We provide two baselines: accuracy when the question-decomposition step is performed by Qwen2.5-7B itself versus a fine-tuned model for decomposition. The fine-tuned model on its own yields no improvements; improvements are due to better decompositions and not from the improvements in general model capabilities from fine-tuning.
Baselines setting
@@
Compute
For our adversarial training experiments in Section 4, we use roughly $100$ hours on a single NVIDIA A100 GPU node. For the result of Section 5, we deployed 8 * NVIDIA H100 GPUs for 40 minutes only to fine-tune the Qwen2.5-7B model on 700 data collected from o3-mini.
Appendix G Decomposition attacks are more effective with jailbreaks
Sometimes, decomposition attacks fail, and the new prompts that are designed to appear benign are actually refused. In these cases, the attacker can apply an additional jailbreak on the refused decomposition(s) in order to obtain a response despite an initial refusal. Using the notation from Table 2, this new protocol corresponds to
[…]
To evaluate this decomposition-then-jailbreak protocol for white-box jailbreaks, we create a new evaluation dataset designed be more solvable for smaller models (Llama3.1 8B) but still challenging (where Qwen2.5 0.5B still struggles). These questions were generated using the same BSD pipeline described in Section 3, but calibrated to provide an appropriate difficulty level for these models (i.e., we used 0.5 as the weak model in the pipeline shown in Figure 2 instead of the more performant 7B model in the Qwen2.5 family of models). We generate 126 easier biology questions with this replacemnt to the pipeline.
As illustrated in Figure 8, the decomposition attack described in Section 5 significantly outperforms GCG attacks, with the latter exhibiting a substantial jailbreak tax [11] that often results in meaningless or incorrect responses. We evaluate the decompose-then-jailbreak approach, where GCG attacks are applied to the decomposed subquestions, and find that it outperforms all other methods (except for the ‘uncensored’ Llama 3.1 8B, which is finetuned to remove refusals and used as a reference for the upper-end of capabilities of the strong model). We fix the total GCG attack steps, ensuring that the steps for all subquestions combined are equivalent to those applied to the original question. This hybrid protocol achieved responses to all decomposed subquestions and increased the misuse rate to 87%, compared to 84% with decompositions alone (and 40% for Llama-3.1 8B).

### v2

@@ F.2 Hyperparameter Details
In Section˜6, we fine-tune Qwen2.5-7B with supervised fine-tuning (SFT) on 700 set of decomposition sub-queries samples collected from o3-mini on MMLU-auxilary, where the number of decompositions varied from 3 to 12 across the samples. The fine-tuning data is highly structured– extensive fine-tuning on such data can lead to overfitting, causing the model to lose its general capabilities. It is known that a fine-tuned model on small or structured datasets memorizes training data rather than overall generalization Lialin et al. (2024). Hence, we found that a small learning rate ($1{\rm e}-6$) with only 2 epochs work the best to circumvent this. We provide an example question and decomposition datapoint below.
MMLU-auxiliary question Question: Every student in Australia must wear their school uniforms when they are at school. There are many kinds of uniforms such as suits, dresses, shorts, skirts, shirts, T-shirts and so on. Class usually begins at 8:$0$ a.m. in most Australian schools, and sometimes it begins at 10:$0$ a.m. You needn’t go to class with your schoolbag, you can put it in your own cabinet and only take a few books. There are many subjects: English, Maths, Biology, Geography, History, Physics, Chemistry, Music, Computer, PE, Chinese Drama and so on. In a computer class, each student sits at a table and does their work. They are very $__$. The teacher goes around them to give directions to them. One class is about two hours. At noon all the students have lunch at school. Their lunch is sandwiches and some fruit. In the afternoon there’s only one class. School ends at 3:30 P.m. Some students can swim or play ball games until 5:30 P.m. Then they must go home. They seldom do homework, even the students in Grade Eleven. This is the school life of the Students in Australia. In a computer class, each student is very $__$. Choices:
@@
Decomp 1)
In the sentence ’In a computer class, each student sits at a table and does their work. They are very $__$.’, what setting is being described, and what are the main actions of the students?
Decomp 2)Based on the sentence ’In a computer class, each student sits at a table and does their work. They are very $__$.’, is the missing adjective more likely to describe the students’ volume level (how much noise they make) or their emotional state? Decomp 3)Which of the following words describes a low level of sound: quiet, noisy, happy, or sad? Decomp 4)What does the phrase ’does their work’ suggest about how much talking or noise the students are making in the computer class context? Decomp 5)In a scenario where students are each working individually on computers and not talking, would you describe the atmosphere as more ’quiet’ or more ’noisy’? Decomp 6)Given the sentence ’Each student sits at a table and does their work. They are very $__$.’ and the four options quiet, noisy, happy, sad, which word logically fills the blank to match the described atmosphere?
F.3 Decomposition Finetuning Ablation
In order to show that the fine-tuning is targeted to help with the quality of the decomposed sub-queries, and does not significantly enhance the general capability of the fine-tuned model, we provide a baseline experiment where the fine-tuned model is used together with Qwen in the decomposition baseline (i.e., fine-tuned model $\to$ Qwen $\to$ Qwen). We compare this with the scenario that Qwen is used for all three parts. As Table˜4 shows, the fine-tuned model does not improve the accuracy when no strong model is deployed.
@@ F.3 Decomposition Finetuning Ablation
Table 4: Distillation leads to misuse because the attacker learns better decomposition strategies, not general gains in capabilities. We provide two baselines: accuracy when the question‑decomposition step is performed by Qwen2.5‑7B itself versus a fine‑tuned model for decomposition. The fine‑tuned model on its own yields no improvements; improvements are due to better decompositions and not from the improvements in general model capabilities from fine-tuning.
F.4 Defense Details
Stateful defense results against decomposition attacks
. In Table˜5, we show the precision and false positive rates for the pointwise and our buffer defense on with of 90% and 99%. As discussed, the buffer defense dramatically outperforms the pointwise defense, maintaining similar precision values even when the signal is very noisy, and the harmful prompts are rare with respect to the benign prompts.
Table 5: Decomposition Attack: Precision and FPR for Pointwise and Buffer Defense
Harmful 90% Recall 99% Recall
Fraction Pointwise Buffer Pointwise Buffer
Prec FPR Prec FPR Prec FPR Prec FPR
1/10 0.156 0.844 0.187 0.813 0.100 0.900 0.187 0.813
1/100 0.015 0.985 0.130 0.870 0.012 0.988 0.119 0.881
1/400 0.004 0.996 0.184 0.816 0.003 0.997 0.172 0.828
1/1000 0.003 0.997 0.139 0.861 0.001 0.999 0.150 0.850
1/4000 0.000 1.000 0.174 0.826 0.000 1.000 0.188 0.812
Naive rolling-window baseline.
Consider a baseline where the entire previous conversation with the user, across all sessions, is put in-context for the monitor model. This baseline corresponds to letting k (the number of previous queries) be the same as m (the buffer size). This works for small values of previous queries (see Figure˜5, for fraction harmful $\frac{1}{2}$, $\frac{1}{4}$, and $\frac{1}{10}$), however our adaptive attacker can overflow this context with benign distractor queries (we assume the context length for the defender is  30 queries– e.g., a realistic defender model may be unlikely to have a context length larger than 32K tokens, which accounts to only [32 benign prompts] x [1K tokens per prompt]). So the defender has to make a design choice: the naive decision would be to select only the latest queries that can fit into the context length of the language model, a “rolling window buffer.” However, the attacker can simply space out their harmful query (the decomposition from a harmful question) such that it is mixed with enough benign queries to overflow the context length. Our buffer gets around this by assuming that the harmful queries will be on average more pointwise harmful than most of their benign counterparts. We find this works reasonably well. Another advantage of our buffer is that it can be cached, and this cache will be refreshed far less than the rolling window buffer. In short, we introduce a naive defense, an adaptive attack, and a less-naive defense, and benchmark them.
Appendix G Decomposition attacks are more effective with jailbreaks
Sometimes, decomposition attacks fail, and the new prompts that are designed to appear benign are actually refused. In these cases, the attacker can apply an additional jailbreak on the refused decomposition(s) in order to obtain a response despite an initial refusal. Using the notation from Table˜2, this new protocol corresponds to
$X\xrightarrow{W}\{X_{i}\}\xrightarrow{S^{*}}\{Y_{i}\}\xrightarrow{W}Y,$ (1)
To evaluate this decomposition-then-jailbreak protocol for white-box jailbreaks, we create a new evaluation dataset designed be more solvable for smaller models (Llama3.1 8B) but still challenging (where Qwen2.5 0.5B still struggles). These questions were generated using the same BSD pipeline described in Section 4, but calibrated to provide an appropriate difficulty level for these models (i.e., we used 0.5 as the weak model in the pipeline shown in Figure˜2 instead of the more performant 7B model in the Qwen2.5 family of models). We generate 126 easier biology questions with this replacement to the pipeline.
(a) Standard decomposition attack.
(b) Decomposition attack with a jailbreak conditional on an initial refusal.
Figure 10: (a) In a standard decomposition attack, a harmful task is broken up into $n$ benign subtasks, which are passed to a strong model. The strong model solutions are put in-context for a weak helpful-only model to help it solve the original harmful task. This attack was first introduced in (Jones et al., 2024, Glukhov et al., 2024). (b) We introduce a decomposition attack variant (Equation˜1). Here, when a benign task is refused, we apply an additional jailbreak. In Figure˜9, we find that the attack is more effective (has higher misuse rate) than the standard decomposition attack or a GCG jailbreak in isolation (we control for number of GCG iterations).
As illustrated in Figure 9, the decomposition attack described in Section 6 significantly outperforms GCG attacks, with the latter exhibiting a substantial jailbreak tax Nikolić et al. (2025) that often results in meaningless or incorrect responses. We evaluate the decompose-then-jailbreak approach, where GCG attacks are applied to the decomposed subquestions, and find that it outperforms all other methods (except for the ‘uncensored’ Llama 3.1 8B, which is finetuned to remove refusals and used as a reference for the upper-end of capabilities of the strong model). We fix the total GCG attack steps, ensuring that the steps for all subquestions combined are equivalent to those applied to the original question. This hybrid protocol achieved responses to all decomposed subquestions and increased the misuse rate to 87%, compared to 84% with decompositions alone (and 40% for Llama-3.1 8B).
Appendix H System prompts
@@
Composer You are tasked with answering a multiple-choice question- please provide the correct answer to the question using your general knowledge, your reasoning, and the information provided in the following sub-questions. You will first receive a list of related questions and answers to those questions that may inform your answer to the main question. Setting: You will need to answer this question: "{question}" Options: {options} Here are some related questions and answers that might inform your answer to the main question: {subtask_responses} Output Format: **Please think step-by-step to reason about the question and provide your answer.** For your final answer, you must provide the letter of the correct answer choice, for example: "ANSWER: <letter>
Appendix I BSD Dataset access request form
We gate access to the BSD dataset due to concerns around misuse, but are committed to allowing access to researchers interested in using the dataset for legitimate reasons. Below, we provide our access request form, available at the link [anonymous].
Thank you for your interest in accessing our Benchmarks for Stateful Defenses (BSD) dataset. As outlined in our access policy, we maintain controlled access to ensure the dataset is used for legitimate safety research while preventing potential harmful applications. To process your request, please provide the following information (feel free to make your answers brief and informal):
Research Purpose:
A description of your intended research objectives and expected outcomes. How you plan to use the BSD dataset specifically.
Research Background:
Brief overview of your research background. Your current institutional affiliation and role.
Technical Details:
What aspects of the dataset are most relevant to your work (misuse uplift measurement, detectability evaluation, etc.). Rough modifications or extensions you intend to make to the evaluation framework (if any).
Data Handling:
Description of your data security measures and storage protocols. Confirmation that you will not redistribute the dataset or derived materials. Please reply to this email with this information, along with any supporting documentation that demonstrates the legitimacy and safety focus of your research. We aim to review all requests promptly.